a fulcrum is placed under the center of a board. in order for two objects to balance on the board, the distance (in feet ) of each object from the center of the board must vary inversely with its weight (in pounds) . what is the distance of each object from the center of the board ?

(7.5 lb)a___________5ft______b (12.5 lb)
the distance between two objects is 5ft

Answer by checkley77(12844)

7.5x=12.5(5-x)
7.5x=62.5-12.5x
7.5x+12.5x=62.5
20x=62.5
x=62.5/20
x=3.125 is the distandce of the 7.5 pound item from the center.
5-3.175=1.875 is the distance of the 12.5 pound weight from the center.
Proof:
7.5*3.125=12.5*1.875
23.4375=23.4375